OpenStudy (anonymous):

I am going to tell you step by step how i did an equation and you can tell me if i got it right

OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -one thirdx + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

so first I have to turn the slope opposite or whatever and I do this by flipping it then making it a positive sense it was negitive adn it would look like this 3/2

jimthompson5910 (jim_thompson5910):

given slope is -1/3 flip the fraction to get -3/1 then flip the sign to get +3/1 or 3/1 or 3

OpenStudy (anonymous):

yeah opps thats what I meant I don't know where the two came from thanks

OpenStudy (anonymous):

Okay so now i put it in point slope formula and this is how it would look...

OpenStudy (anonymous):

\[Y -3 = 3(x +1)\] then I do the distributive property on the right side \[Y - 3= 3x +3\] Then I add the three from the left to both sides and it would look like this \[y = 3x \] so did i get it?

OpenStudy (anonymous):

y = 3x +6

jimthompson5910 (jim_thompson5910):

not sure how you got y = 3x, but y = 3x + 6 is correct

OpenStudy (anonymous):

no I fixed it lol

jimthompson5910 (jim_thompson5910):

everything else is correct though, so very nice work

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

sure thing